# Variety and Quality over Quantity: Towards Versatile Instruction Curation

### 12 Dec 2023 v1

@@ Abstract
Instruction fine-tuning, involving the refinement of pre-trained LLMs using datasets accompanied by natural instructions, is a powerful approach. However, its effectiveness is hindered by the redundancy and deficiencies in LLM-generated instruction datasets. In this paper, we introduce a highly effective and versatile paradigm for selecting diverse and high-quality instruction-following data from fine-tuning datasets. We first employ the dataset enhancement and expansion to augment the dataset with more diverse and high-quality data, then we apply variety compression and quality compression sequentially to curate the desired dataset. Our experimental results showcase that, even with a limited quantity of high-quality instruction data, LLMs consistently maintain robust performance across both natural language understanding tasks and code generation tasks. Notably, they outperform models trained on significantly larger instruction datasets in certain instances.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have gained prominence for their remarkable effectiveness in natural language comprehension tasks (OpenAI, 2023; Yang et al., 2023). High-quality pretrained LLMs like LLaMA2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023) and StarCoder (Li et al., 2023b) are readily available, facilitating their customization for versatile applications. One popular fine-tuning approach, known as instruction tuning (Wei et al., 2022; Ouyang et al., 2022), involves fine-tuning pre-trained LLMs using datasets accompanied by natural language instructions. Its relative simplicity and affordability make it a preferred method for improving language models.
Figure 1: Proposed Instruction Dataset Curation Paradigm Overview
However, a significant challenge in the instruction tuning process arises from the observation that fine-tuning with larger instruction datasets does not always guarantee better results, yet demanding more computational resources. This issue stems from the redundancy and deficiencies in LLM-generated instruction datasets (Luo et al., 2023). The LLM-generated dataset, proposed as a substitute for resource-intensive manual data selection, lacks supervision regarding the diversity and quality of the generated data. Dataset redundancy poses the risk of the model overfitting to the fine-tuning data, thereby hindering its capacity for generalization. Concurrently, the inclusion of low-quality or potentially harmful data can degrade the performance of LLMs.
The concern surrounding the instruction fine-tuning process has prompted researchers to explore methods for fine-tuning large models using a limited quantity of high-quality data in order to attain robust performance. Encouragingly, several recent studies have validated the efficacy of this approach, demonstrate that large language models can achieve impressive results even when provided with a limited quantity of high-quality instruction-following data (Zhou et al., 2023; Wei et al., 2023). However, current research is still primarily focused on limited models and tasks. There is still a lack of systematic exploration on compression principles and techniques.
Inspired by the previous work on curating fine-tuning dataset for LLMs, we propose a new LLM instruction dataset curation paradigm, presented in figure 1. This paradigm is designed to generate a diverse and high-quality dataset for fine-tuning LLMs, while significantly reduce the quantity of the dataset. Our paradigm is composed of three steps. First, we employ the "Dataset Enhancement and Expansion", enhancing the accuracy of the original instructions’ answers and introduce new high-quality data. Subsequently, the "Variety Compress" approach eliminates redundancy from both the original and newly generated data. Finally, the "Quality Compress" process assesses the quality of each item in the compressed data, selecting high-quality ones to form the final curated dataset.
@@
To validate the effectiveness of our paradigm, we employ the final curated instruction datasets for fine-tuning open-source LLMs. We conduct extensive experiments to evaluate these fine-tuned LLMs’ performance in both natural language understanding (NLU) tasks and code generation tasks, considering 6 metrics. Results consistently show that LLMs maintain robust performance even with a limited quantity of high-quality instruction data. In certain cases, they even surpass models trained on larger datasets.
Our contribution can be summarized as follows:
1. We propose a highly effective and versatile paradigm for the selection of diverse and high-quality instruction-following data from fine-tuning datasets.
2. Our experiment results demonstrate that with a small quantity of instruction data selected by our paradigm, LLMs can still maintain robust performance, and in some cases, even outperform models trained with larger instruction datasets.
3. Our findings provide valuable insights, suggesting that instead of indiscriminately feeding large volumes of data into LLMs during training, selecting a smaller quantity of high-quality, multifaceted data for fine-tuning not only conserves time and resources but also preserves and, in some tasks, enhances performance.
@@ 2 Related Works
2.1 Instruction Tuning
While LLMs have demonstrated cross-task generalization capabilities, they often fall short in aligning with users’ prompts when applied to real-world user communications. To overcome this limitation, OpenAI enlisted professional human annotators to generate an extensive corpus with human instructions. Leveraging this instruction dataset, OpenAI fine-tuned its GPT-3 model to create InstructGPT (Ouyang et al., 2022). Subsequently, recent research has pivoted towards automatic instruction generation frameworks. Alpaca (Taori et al., 2023) proposed adopting the self-instruct method by utilizing ChatGPT to generate data for fine-tuning. Vicuna (Chiang et al., 2023) utilized data collected from ShareGPT.com, a platform where users share their conversations with ChatGPT, for fine-tuning their models. WizardLM (Xu et al., 2023) and WizardCoder (Luo et al., 2023) introduced the Evol-Instruct method, which involves evolving existing instruction data to generate more diverse and complex data.
LLM-generated datasets often fall short in dataset variety and quality, due to limited measures to guarantee and control the output of LLMs. This hidden issue further constrains the potential for performance improvement in LLMs.
@@ 2.2 Instruction Dataset Curation
To enhance the model’s performance following instruction fine-tuning, some researchers have concentrated on filtering out low-quality data during the fine-tuning stage. LIMA (Zhou et al., 2023) demonstrates that fine-tuning a robust pre-trained language model on 1000 high-quality, human-curated examples can yield remarkable and competitive results. Instruction Mining (Cao et al., 2023) introduces a linear rule for selecting high-quality instruction data, eliminating the need for human annotation, as seen in LIMA. Du et al. present a model-oriented data selection (MoDS) approach, which selects instruction data based on a new criteria considering three aspects: quality, coverage and necessity. Li et al. introduce a self-guided methodology for LLMs to autonomously discern and select cherry samples from vast open-source datasets, effectively minimizing manual curation and potential cost.
@@
3 Methodology
As shown in figure 1, our paradigm consists of three steps: Data Enhancement and Expansion, Variety Compress and Quality Compress. In this section, we’ll introduce these three steps in details.
3.1 Data Enhancement and Expansion
The original dataset often contains brief and straightforward prompts, resulting in simplistic responses lacking significant semantic information. These instances of low-quality data are part of the original dataset, suggesting that relying solely on compressing it would yield limited performance improvements. To boost the quality of our compressed dataset, we find data enhancement and expansion necessary. The workflow of these two phase is presented in figure 2.
Figure 2: Instruction Dataset Enhancement and Expansion
3.1.1 Data Enhancement
In data enhancement phase, we replace the original dataset’s output section with GPT-4’s response. The output section of the original instruction dataset is often generated by humans or other LLMs. While generally accurate and aligned with users’ intentions, these outputs frequently lack comprehensive explanations due to brevity. For instance, a prompt like "Write a function to generate the nth Fibonacci number" might yield a concise Python code answer in the original dataset, lacking detailed coding logic and considerations for boundary cases.
Recognizing GPT-4’s exceptional natural language and code comprehension, we choose to utilize GPT-4 to re-generate the "output" section in the dataset. The prompt given to GPT-4 comprises the original dataset’s "instruction" and "input" sections. Compared to the original output, GPT-4 tends to generate more detailed answers, including comments for programming requests and additional reasoning steps to address logic questions.
3.1.2 Data Expansion
Inspired by the instruction rewrite method proposed in WizardLM and WizardCoder (Xu et al., 2023; Luo et al., 2023), our approach is similarly dedicated to producing instructions that are diverse and intricate. We implement the "Instruction-Elevate" process, where we instruct GPT-4 to function as a prompt re-writer, generating challenging instructions based on specified generation rules.
Given the variation in instruction dataset content for Natural Language Understanding (NLU) and code generation tasks, we provide distinct settings for GPT prompts to increase complexity. Additional details are available in Appendix A.
After generating new instructions and inputs, we use GPT-4 to produce corresponding outputs, concluding an expansion round. Each new instruction item corresponds to an original one, resulting in a dataset of the same size as the original after each round. We iterate this process for $k$ rounds, combining the expanded datasets with the enhanced dataset to create a final expanded dataset.
3.2 Variety Compress
The "Variety Compress" step aims to eliminate duplicated data from both the original dataset and the expanded dataset, generated in the previous step. Duplicated items may lead LLMs to overfit on specific topics, reducing their ability to generalize knowledge. The workflow is depicted in Figure 3. To identify duplication, we opt to analyze embeddings of the expanded dataset.
Initially, GPT generates embeddings with 1536 dimensions for each item, but this proves unwieldy for analysis. To address this, we aim to reduce the embedding dimension and devise a method to represent data differentiation.
Figure 3: Variety Compress
3.2.1 Normalization & Dimension Reduction
We first calculate the covariance matrix of the given features, perform eigenvalue decomposition on the covariance matrix to obtain eigenvalues and eigenvectors, then choose top $k$ eigenvectors corresponding to the largest $k$ eigenvalues, where $k$ is the target reduced dimension. It’s essential to acknowledge that as we compress the dimension of feature vectors, some information loss is inevitable. Careful consideration is needed when selecting the target dimension denoted as $k$ to ensure retention of the majority of the original information.
Before directly reducing the embedding data’s dimension, feature normalization is necessary. Given covariance matrix’s sensitivity to varying feature scales, normalization becomes crucial as larger-scale features may dominate the covariance matrix, impacting the computation of principal components and potentially introducing bias. Normalization mitigates this bias, ensuring equal contribution from each feature. After normalization, we undertake the following steps to ascertain the target reduced dimension $k$. We designate $k$ as the reduced dimension, aiming for a value where $k$ features can retain approximately 95% of the variance present in the original embeddings, thus preserving a substantial amount of information.
3.2.2 Row Variance for Dimension-reduced Features
Following dimension reduction, data undergoes linear transformation and projects into a new coordinate system defined by selected $k$ features. Reduced-dimensional data corresponds to this projection, allowing analysis of row variance to identify items with significant differences. Row variance measures variability in the reduced space across all $k$ dimensions for each data point. High variance suggests substantial positional changes along that dimension, aiding in identifying diverse data points. After calculating row variances, we select items with the highest 20% row variances to construct the variety-compressed dataset, ensuring diverse instructions are chosen.
3.3 Quality Compress
Figure 4: Quality Compress
"Quality Compress" is designed to identify high-quality instruction data from a compressed dataset. High-quality instructions and their outputs should meet specific criteria: clear and detailed instructions, presenting a reasonable level of complexity for advanced LLMs. The expected output should align accurately with instructions and provide comprehensive explanations.
Rating instruction data quality is challenging without official quantitative metrics. Due to the dataset’s size and associated costs, enlisting professional annotators for individual scoring is unrealistic. Therefore, we leverage GPT-4, an advanced LLM, as an instruction scorer, generating GPT quality score for each instruction. Furthermore, recognizing that the total length of instruction data indicates information richness, we incorporate this into the quality score calculation. Unlike other metrics, length is quantitatively analyzable. A positively correlated mapping function derives a lengthwise semantic score based on the instruction data’s length. Combining the GPT quality score and lengthwise semantic score produces the final quality score. Instruction items with high-quality scores compose the final quality-compressed dataset, as illustrated in Figure 4.
3.3.1 GPT Quality Evaluation
We engage GPT-4 to evaluate the instruction data across four dimensions: accuracy, explanation, clarity, and difficulty, each contributing to a distinct portion of the total GPT score.
Accuracy (35%): Reflects the correctness of the output solution concerning the instruction’s problem. A high score is awarded for a solution that perfectly addresses all requirements.
Explanation (25%): Applied to the output, this metric rewards detailed explanations, such as comments in coding tasks and reasoning steps within a chain of thoughts.
Clarity (15%). Assesses how well the instruction articulates requests, emphasizing clear details and avoiding ambiguity or overly general questions.
Difficulty (25%). Focuses on the complexity of the instruction’s problem. Higher scores are given for requests with reasoning constraints and additional depth of thought.
These metrics’ proportions are assigned based on their contributions to overall instruction quality. The template guiding the GPT-4 scores is presented in Appendix B.
@@ 3.3.1 GPT Quality Evaluation
In our practical experience, we observed that GPT-4 consistently assigns high total scores (over 90) to all instruction data, even when explicitly instructed not to do so. This behavior stems from GPT-4’s training as an overall friendly and positive agent, posing a challenge by reducing differentiation in GPT quality scores. To address this, we implement the following steps for more reasonable and differentiated scores:
@@
3.3.2 Lengthwise Semantic Evaluation
The lengthwise semantic complexity score is determined by a linear mapping function. Instructions with the shortest total length, including both the instruction and output, receive the lowest score, while those surpassing a specified threshold attain the highest score.
Combining the GPT quality assessment scores and lengthwise semantic scores produces the final quality score, facilitating the selection of high-quality items. In Appendix C, we present an overview of the total quality score distributions across both tasks. Our meticulous efforts ensure substantial differentiation in quality scores, allowing precise identification and selection of high-quality data.

### 27 Dec 2023 v2

# Rethinking the Instruction Quality: L IFT is What You Need

@@ Abstract
Instruction tuning, a specialized technique to enhance large language model (LLM) performance via instruction datasets, relies heavily on the quality of employed data. Existing quality improvement methods alter instruction data through dataset expansion or curation. However, the expansion method risks data redundancy, potentially compromising LLM performance, while the curation approach confines the LLM’s potential to the original dataset. Our aim is to surpass the original data quality without encountering these shortcomings. To achieve this, we propose LIFT (LLM Instruction Fusion Transfer), a novel and versatile paradigm designed to elevate the instruction quality to new heights. LIFT strategically broadens data distribution to encompass more high-quality subspaces and eliminates redundancy, concentrating on high-quality segments across overall data subspaces. Experimental results demonstrate that, even with a limited quantity of high-quality instruction data selected by our paradigm, LLMs not only consistently uphold robust performance across various tasks but also surpass some state-of-the-art results, highlighting the significant improvement in instruction quality achieved by our paradigm.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have gained prominence for their remarkable effectiveness in natural language comprehension tasks (OpenAI, 2023; Yang et al., 2023). High-quality pretrained LLMs are readily available, facilitating their customization for versatile applications. One popular fine-tuning approach, known as instruction tuning (Wei et al., 2022; Ouyang et al., 2022), involves fine-tuning pre-trained LLMs using datasets accompanied by natural language instructions. Its relative simplicity and affordability make it a preferred method for improving LLMs’ performance on specific tasks.
Figure 1: Instruction Dataset Curation Paradigm LIFT
The effectiveness of instruction tuning heavily relies on the quality of the instruction dataset. Currently, two prevalent methods for constructing instruction datasets are manual crafting (Ouyang et al., 2022) and LLM-generated (Chiang et al., 2023; Taori et al., 2023). However, these methods often fall short in terms of quality. Human-crafted datasets depend on human annotators to generate a substantial corpus with human instructions, leading to the instruction dataset lacking detailed context and explanation. It may also contain vague or objective descriptions. On the other hand, LLM-generated methods use advanced LLMs to generate or complete instructions and responses but lack supervision regarding the diversity and quality of the generated data.
The concern regarding instruction datasets’ quality has prompted researchers to explore methods for improving their quality. Current approaches to quality enhancement can be broadly categorized into two groups: data expansion and data curation. Data expansion methods involve leveraging advanced LLMs, such as GPT-4, with a suitable prompt template to generate new instructions and corresponding answers based on the original dataset (Xu et al., 2023; Luo et al., 2023; Taori et al., 2023). On the other hand, data curation methods involve the meticulous selection of high-quality data from the original dataset based on specific quality evaluation criteria (Zhou et al., 2023; Du et al., 2023). We posit that both methods essentially function as data distribution transfers: expansion enables the distribution to cover a broader range of data subspaces, typically characterized by higher quality, while curation concentrates the distribution on a higher-quality subset of the original dataset.
However, both current methods have limitations that impede their capacity to further enhance performance. Expansion methods introduce redundancy into the dataset as the newly generated instructions are typically based on the original ones. The effectiveness of curation methods heavily relies on the quality of the original dataset, limiting the quality of the curated dataset. These limitations force these methods to rely on specific expansion or curation strategies to achieve better performance on certain benchmarks, at the expense of losing the ability to generalize the approach.
@@
In this paper, we delve deeper into the quality distribution transfer process to address the mentioned problems. We propose a novel paradigm for improving LLM instruction quality, termed LIFT (LLM Instruction Fusion Transfer). LIFT is designed to amalgamate the advantages of data expansion and curation, mitigating their shortcomings to generate a diverse and high-quality dataset while significantly reducing quantity. Our paradigm consists of two phases. Firstly, we employ "Dataset Distribution Expansion", broadening the data distribution to cover more high-quality subspaces. Then, we utilize "Dataset Variety and Quality Curation" to eliminate redundancy and densify the data distribution, focusing on the high-quality segments of overall data subspaces. To validate the effectiveness of LIFT, we employ the finally curated instructions for fine-tuning open-source LLMs. Through extensive experiments evaluating the performance of these fine-tuned LLMs in both natural language understanding (NLU) tasks and code generation tasks, the results consistently demonstrate that LLMs maintain robust performance even with a limited quantity of high-quality instruction data. In certain cases, they even outperform models trained on larger datasets. To summarize, our main contribution are:
  • We propose a highly effective and versatile paradigm, LIFT, which challenges the conventional single-mode enhancement for instruction datasets. LIFT rethinks data quality by focusing on data distribution transfer. It aims to elevate the quality of the instruction dataset to new heights, overcoming redundancy and quality limitations present in current methods.
  • Throughout the expansion and curation phases of the paradigm, we prioritize both variety and quality as essential goals for enhancement. Unlike existing works that concentrate on only one stage, we posit that these characteristics play a crucial role in bringing more high-quality data into the final curated dataset.
  • Our extensive experimental results demonstrate that, with a significantly reduced quantity of high-quality instruction data selected by our paradigm, LLMs consistently achieve nearly-SOTA or SOTA performance across various benchmarks. This provides valuable insights, suggesting that a selective approach based on the principles of data distribution transfer is not only more effective but also cost-effective compared to the indiscriminate feeding of large volumes of data into LLMs during training.
@@ 2 Related Works
2.1 Instruction Dataset Expansion
@@ 2.2 Instruction Dataset Curation
To enhance the model’s performance following instruction fine-tuning, some researchers have concentrated on filtering out low-quality data during the fine-tuning stage. LIMA (Zhou et al., 2023) demonstrates that fine-tuning a robust pre-trained language model on 1000 high-quality, human-curated examples can yield remarkable and competitive results. Instruction Mining (Cao et al., 2023) introduces a linear rule for selecting high-quality instruction data, eliminating the need for human annotation, as seen in LIMA. Du et al. (2023) present a model-oriented data selection (MoDS) approach, which selects instruction data based on a new criteria considering three aspects: quality, coverage and necessity. Li et al. (2023a) introduce a self-guided methodology for LLMs to autonomously discern and select cherry samples from vast open-source datasets, effectively minimizing manual curation and potential cost.
@@
3 LLM Instruction Fusion Transfer
3.1 Data Distribution Transfer
Current methods for enhancing instruction quality, whether through data expansion or curation, do enhance the original dataset to some extent, thereby improving the performance of LLMs on specific tasks when using these enhanced datasets for instruction tuning. However, the effectiveness of these methods is constrained by inherent limitations. To scrutinize these limitations and explore innovative approaches to break from conventional enhancement modes, we propose a novel perspective for rethinking instruction data quality: data distribution transfer.
3.1.1 Analysis of Existing Methods
Our hypothesis is that, during the process of enhancing instruction quality, there is a transfer of data distribution from the original dataset to the final enhanced dataset. This transfer increases the quantity or proportion of high-quality data. In data expansion methods, generating high-quality instructions based on the original ones effectively extends the coverage of the high-quality data subspace within the original data distribution, thereby increasing the quantity of high-quality data in the final distribution. On the other hand, in data curation methods, by using carefully designed quality evaluation metrics, low-quality components are removed from the final distribution, directing the distribution to concentrate on high-quality data and increasing its proportion in the final distribution.
From this perspective, we can easily identify the limitations in these methods. The areas around the original instructions may contain similar ones, leading to redundancy in the final distribution. Moreover, low-quality instructions and those derived from them persist in the final distribution, maintaining a proportion similar to the original dataset. On the other hand, curation methods select a portion of high-quality instructions from the original dataset, resulting in a decrease in the total number of high-quality instructions. If the original dataset has a limited number of high-quality instructions, the quality of the curated dataset will significantly decrease.
3.1.2 Fusing Expansion and Curation
Analyzing the data distribution transfer patterns of expansion and curation, we suggest that their integration can effectively address their individual limitations. Data expansion broadens subspaces, enabling the curation method to explore beyond the original distribution. Conversely, data curation assists in identifying duplicates and low-quality items from the expansion, contributing to a more concentrated and refined distribution.
Building on these insights, we introduce a novel paradigm called LIFT (LLM Instruction Fusion Transfer). Comprising two phases, this paradigm orchestrates the distribution of instruction data as follows: in the "Dataset Distribution Expansion" phase, we broaden the data distribution to encompass more diverse and high-quality subspaces, acknowledging the presence of duplications at this stage. Subsequently, in the "Dataset Variety and Quality Curation" phase, we systematically eliminate redundancy and low-quality elements, creating a densified distribution for the final curated dataset. These two phases are intricately connected, ensuring a smooth transfer of data from the original dataset to the final curated dataset.
3.2 Paradigm LIFT
As mentioned in the above section, our paradigm LIFT follows a two-stage structure. In both stages, we value the diversity and quality as the crucial criterion since we believe the "Dataset Distribution Expansion" and "Dataset Variety and Quality Curation" equally contribute to the quality enhancement.
3.2.1 Dataset Distribution Expansion
The goal of dataset distribution expansion is to encompass a more diverse and high-quality range of data within the distribution, while ensuring a certain distance from the original instructions. To achieve this, it is crucial to employ carefully designed instruction-generation prompts. Drawing inspiration from the instruction rewrite method proposed by Xu et al., our approach focuses on generating diverse and intricate instructions. We guide GPT-4 to act as a prompt rewriter, generating challenging instructions based on specified generation rules. Considering the variation in content for NLU and code generation tasks within the instruction dataset, we configure distinct settings for GPT prompts to enhance complexity. For further details, refer to Appendix A. We iterate this process for $k$ rounds, merging the expanded datasets with the original dataset to create the final expanded dataset.
3.2.2 Dataset Variety and Quality Curation
Figure 2: Variety Compress with Dimension Reduction and Row Variances
Figure 3: Quality Compress with GPT Quality Evaluation and Lengthwise Semantic Evaluation
An effective curation method ought to eliminate duplicated or low-quality instructions from the original dataset, while preserving representative and high-quality ones. To meet this criterion, curation should be approached with meticulous attention to both variety and quality.
Current variety curation typically involves clustering methods such as k-means or spectral clustering, which initially segment the original data distribution into several small groups, followed by the selection of representative items from these groups. We argue that this approach may lack generalizability and be less effective when dealing with new datasets. This is because these methods require prior knowledge of the number of clusters, and choosing cluster numbers that are either too large or too small may reduce their effectiveness in selecting representatives.
Our variety curation method take another route, as depicted in Fig. 2. Initially, GPT generates embeddings with 1536 dimensions for each item, which proves unwieldy for analysis. To address this, we aim to reduce the embedding dimension and devise a method to represent data differentiation. We achieve this by calculating the covariance matrix of the given features and performing eigenvalue decomposition on the covariance matrix to obtain eigenvalues and eigenvectors. We then choose the top $k$ eigenvectors corresponding to the largest $k$ eigenvalues, where $k$ is the target reduced dimension. This process allows us to analyze row variance on the dimension-reduced features to identify items with significant differences. Row variance measures variability in the reduced space, and high variance suggests substantial positional changes along that dimension, aiding in identifying diverse data points. We select items with the highest 20% row variances to construct the variety-curated dataset. This method doesn’t require any prior statistical knowledge of the dataset, making it versatile and effective for all tasks.
Following variety curation is the quality curation phase, where we discern high-quality instruction data. Rating the quality of instruction data is challenging due to the lack of official quantitative metrics. Given the dataset’s size and associated costs, employing professional annotators for individual scoring is impractical. Therefore, we turn to GPT-4 as an instruction scorer, generating GPT quality scores for each instruction. GPT-4 evaluates the instruction data across four dimensions: accuracy, explanation, clarity, and difficulty. The proportions assigned to these metrics are based on their contributions to overall instruction quality. The template guiding the GPT-4 scores is presented in Appx. B.
@@ 3.2.2 Dataset Variety and Quality Curation
In our practical experience, we observed that GPT-4 consistently assigns high total scores to all instruction data, even when explicitly instructed not to do so, posing a challenge by reducing differentiation in GPT quality scores. To address this, we implement the following steps for more reasonable and differentiated scores:
@@
Furthermore, recognizing that the total length of instruction data indicates information richness, we incorporate this into the quality score calculation, too. Unlike other metrics, length is quantitatively analyzable. A positively correlated mapping function derives a lengthwise semantic score based on the instruction data’s length. Combining the GPT quality score and lengthwise semantic score produces the final quality score. Instruction items with high-quality scores compose the final quality-compressed dataset, as illustrated in Fig. 3. In Appx. C, we present an overview of the total quality score distributions across both tasks. Our meticulous efforts ensure substantial differentiation in quality scores, allowing precise identification and selection of high-quality data.